OpenStudy (anonymous):

Determine whether 2 is a zero of the polynomial P(x) = 4x3 – 5x2 + 3x – 10 by using the Remainder Theorem. Explain how the Remainder Theorem is useful in finding the zeros of a polynomial function.

OpenStudy (anonymous):

Use the caret ^ (above the 6) to denote an exponent. P(x) = 4x^3 - 5x^2 + 3x - 10 P(2) = 32 - 20 + 6 - 10 = 8 2 is NOT a zero of the polynomial...the remainder is NOT zero !